OpenStudy (eva12):

1) Parameterize the plane that contains the three points (2,3,−4), (2,−10,−4), and (40,5,5). r⃗ (s,t)=

OpenStudy (anonymous):

c^2=x^2+y^2+z^2

OpenStudy (anonymous):

Let P=(2,3,−4), Q= (2,−10,−4), and R=(40,5,5) compute the vectors PQ=Q-P=(0,-13,0) PR=R-P=(38,2,9)

OpenStudy (anonymous):

One parameterization would be r(s,t) =P + s PQ + t PR= (2 + 38 t, 3 - 13 s + 2 t, -4 + 9 t)

OpenStudy (anonymous):

notice that r(0,0)=P r(0,1)=R r(1,0)= Q